A rectangular box with a square base contains 8 cubic feet. if the height is 6 inches, how many feet are there in each side of the base?
Answer by ankor@dixie-net.com(22740)

A rectangular box with a square base contains 8 cubic feet. if the height is 6 inches, how many feet are there in each side of the base?
:
Let x = side of the base (in ft)
:
It's square, therefore
x^2 = area of the base
:
Given the height as 6 inches, which is .5 ft therefore:
.5x^2 = volume of the box
:
Given the volume as 8 cu/ft, therefore:
.5x^2 = 8
x^2 = 16
x = Sqrt(16)
x = 4 ft is the length of the side
:
:
Check by finding the cu ft:
4 * 4 * .5 = 8